OpenStudy (anonymous):

A cameraman on a pickup truck is traveling westward at 20 km/h while he videotapes a cheetah that is moving westward 30 km/h faster than the truck. Suddenly, the cheetah stops, turns, and then runs at 45 km/h eastward, as measured by a suddenly nervous crew member who stands alongside the cheetah's path. The change in the animal's velocity takes 2.0 s. What are the (a) magnitude and (b) direction of the animal's acceleration according to the cameraman and the (c) magnitude and (d) direction according to the nervous crew member?

OpenStudy (vishal_kothari):

Acceleration from the perspective of the nervous crew member a = (Vf - Vi)/t Choose eastward as negative and westward as positive: a = (-45 km/h - 30 km/h)/2.0s a = (- 37.5 km/h)/s or (- 37.5 km/h)/s x (1000 m/1km) x (1h/3600 s) = - 10.42 m/s^2